# Distribution estimation via Flow Matching with Lipschitz guarantees

### v1

@@
(Karlsruhe Institute of Technology)
@@ 1 Introduction
Flow Matching, introduced by Lipman et al. (2023), has recently attracted attention for its simplicity compared to diffusion models (Sohl-Dickstein et al., 2015), which have long been regarded as state-of-the-art. Unlike generative adversarial networks (Goodfellow et al., 2014), which learn $\psi$ directly, Flow Matching learns a time-depended vector field $v_{t}$ that transports mass from $\mathbb{U}$ to $\mathbb{P}^{*}$ via the ODE for $t\in[0,1]$:
@@
Earlier, normalizing flows (Tabak and Vanden-Eijnden, 2010; Tabak and Turner, 2013), popularized by Rezende and Mohamed (2015) and Dinh et al. (2015) used discrete invertible mappings with the change-of-variables formula. Continuous normalizing flows (Chen et al., 2018) extended this to continuous mappings using neural ODEs trained via maximum likelihood, which circumvented the calculation of the determinant of the Jacobian. However, CNFs still require significant computational costs due to the necessity of ODE simulations (Grathwohl et al., 2019).
Flow Matching avoids these simulations by using a simple least-squares loss
@@
where $v_{t}(\cdot|Y)$ and $p_{t}(\cdot|Y)$ are known by design. In some settings, this resembles the score matching objective (Hyvärinen, 2005; Vincent, 2011) used in diffusion models (Song et al., 2020). Given $n$ i.i.d. observations from $\mathbb{P}^{*}$, minimizing the empirical counterpart of (2) over a set of functions $\mathcal{M}$ and solving the ODE (1) leads to the estimator $\hat{\psi}$. The statistical question is how well $\mathbb{P}^{\psi_{1}(Z)}$ approximates $\mathbb{P}^{*}$ depending on the number of observations $n$ and the function class $\mathcal{M}$.
@@
Algorithms based on Flow Matching have been successfully applied in a range of domains, including text-to-speech (Guo et al., 2024), text-to-image generation (Yang et al., 2024; Esser et al., 2024), molecular and protein structure design (Dunn and Koes, 2024; Bose et al., 2024), and surrogate modeling in high-energy physics (Bieringer et al., 2024). The method has also been extended to various theoretical settings: Atanackovic et al. (2024) adapted it to interacting particle systems, Gat et al. (2024) explored discrete spaces, Chen and Lipman (2024) generalized it to Riemannian manifolds, and Kerrigan et al. (2023) extended it to function spaces.
@@
The analysis of Flow Matching is more fragile compared to Diffusions, as Girsanov’s theorem does not apply, thus making the strategy used in Chen et al. (2023c) inapplicable. Using Grönwall’s Lemma, the natural and standard approach for stability estimates of ODEs, results in exponential dependency on the Lipschitz constant of the approximated vector field. Consequently, the statistical results so far are rather limited: Fukumizu et al. (2024) use a slight adaptation and introduce stopping times to transfer some results from diffusion models but they do not circumvent the growth of the Lipschitz constant, which causes difficulties in their proof. Gao et al. (2024b) use the same adaptation and obtain rates in the Wasserstein 2 distance for certain types of unknown distributions. Stéphanovitch et al. (2025) study deterministic ODE sampling, but, due to their general setting, examine the infinite-time setting of Diffusions. Kunkel and Trabs (2025b) showed that Flow Matching attains minimax optimal rates, but they use an over parameterized setting which does not align with practical implementations. Due to the excess of parameters, they are able to compensate the rising Lipschitz constant.
@@
In this work, we first focus on a detailed study of the Lipschitz constant of the intrinsic "true" vector field of Flow Matching. By providing upper and lower bounds, we demonstrate that the choice of the variance function and the behavior of covariance of the reweighted unknown distribution is crucial. This function governs the level of smoothing depending on $t\in[0,1]$. While current results focus on fixed choices of variance functions, we study the Lipschitz constant in terms of this function. We derive assumptions on the unknown distribution that lead to control over the Lipschitz constant for various types of variance functions. Our general assumptions are validated for a class of distributions with unbounded support considered by Stéphanovitch et al. (2025).
@@
Subsequently, we employ Bernstein-type bounds to derive an oracle inequality for estimating the vector field. However, we carefully adapt these classical methods to address the theoretical challenges of Flow Matching. When bounding the error of the Flow Matching distribution estimator in the Wasserstein $1$ distance, standard approaches from the analysis of diffusion models usually increase the negative impact of the dimension on the rate. This issue can be circumvented up to a small error by exploiting the smoothness of the vector field. Using this approach allows us to build on existing results on network approximation. Our focus is on feedforward rectified linear unit (ReLU) networks, hence we do not require smooth networks for our results. ReLU nets with logarithmically increasing depth and a polynomial number of non-zero weights suffice to achieve improved convergence rates. Additionally, we exploit the density path of the true vector field to take advantage of the smoothness of the unknown distribution. Thus, our rates are faster in high dimensions than those previously obtained in Gao et al. (2024b), while allowing for considerably smaller networks than in Kunkel and Trabs (2025b).
@@
As Lipman et al. (2023) noted, Flow Matching is closely connected to diffusion models. For a comprehensive survey of generative diffusion models, refer to the work of Cao et al. (2024). Even when the model is not built to be consistent with diffusion models, approximating a score function is similar to approximating the vector field in Flow Matching. The statistical analysis of score matching and diffusion models are an active area of research, see for instance Chen et al. (2023a), Chen et al. (2023b), Chen et al. (2023c), Oko et al. (2023), Tang and Yang (2024), Azangulov et al. (2024), Zhang et al. (2024) , Yakovlev and Puchkin (2025), Stéphanovitch et al. (2025).
@@
Section˜2 introduces the problem setting mathematically and recalls all the definitions needed to define the conditional Flow Matching objective from Lipman et al. (2023). Section˜3 then discusses the use of the Wasserstein distance and shows a classical first error decomposition using the aforementioned Grönwall’s lemma. After that, we introduce the model based on the choice $\mathbb{U}=\mathcal{N}(0,I_{d})$, which will be the focus of the subsequent work. In Section˜4 we establish upper and lower bounds of the Lipschitz constant. We then show that bounding the Lipschitz constant requires careful investigation of the covariance of the reweighted unknown distribution. Next, we demonstrate that, under certain assumptions, we can construct Flow Matching such that the Lipschitz constant is bounded. Following that, we present a class of densities that satisfies our assumptions. In Section˜5 we derive an oracle inequality and finally obtain a rate of convergence using feedforward ReLU networks. All proofs are deferred to Section˜7.
@@ 1.1 Some notations
We use the notation $X\sim p$ as shorthand for $X\sim\mathbb{P}$ when the distribution $\mathbb{P}$ of the random variable $X$ admits a density $p$ with respect to the Lebesgue measure.
The symbol $\lesssim$ denotes an inequality of the form $a\leq c\cdot b$, where $a,b\in\mathbb{R}$ and $c$ is a constant independent of the sample size $n$. In the same setting, the symbol $a\sim b$ denotes equality up to a constant. For $d\in\mathbb{N}$, we equip $\mathbb{R}^{d}$ with the standard Euclidean norm $|\cdot|$. For a matrix $A\in\mathbb{R}^{d\times d}$, we denote the spectral norm by $\|A\|$. Given a function $f\colon\Omega\rightarrow\mathbb{R}^{d^{\prime\prime}}$, where $\Omega\subset$ $\mathbb{R}^{d^{\prime}}$, the supremum norm and the $L_{1}$-norm are defined by
$\|f\|_{\infty,\Omega}\coloneqq\sup_{x\in\Omega}|f(x)|\quad\text{and}\quad\|f\|_{1,\Omega}\coloneqq\int_{\Omega}|f(x)|\;\mathrm{d}x.$
When $\Omega=\mathbb{R}^{d^{\prime}}$, we abbreviate $\|f\|_{\infty,\mathbb{R}^{d^{\prime}}}$ as $\|f\|_{\infty}$, and similarly $\|f\|_{1,\mathbb{R}^{d^{\prime}}}=\|f\|_{1}$. For $d^{\prime\prime}=1$ and $\alpha\in(0,1]$, the Besov space $B_{1,\infty}^{\alpha}(\Omega)$ is defined as
$B_{1,\infty}^{\alpha}(\Omega)\coloneqq\big{\{}f\in L_{1}(\Omega)\colon|f|_{B_{1,\infty}^{\alpha}(\Omega)}\coloneqq\sup_{t>0}t^{-\alpha}\omega_{1}(f,t)_{1}<\infty\big{\}},$
where the modulus of smoothness is given by
$\omega_{1}(f,t)_{1}\coloneqq\sup_{0<|h|\leq t}\int|f(x)-f(x+h)|\;\mathrm{d}x,\quad t>0,x\in\Omega.$
If $x+h\notin\Omega$, the integrand is taken to be zero. The norm on $B_{1,\infty}^{\alpha}(\Omega)$ is defined by
$\|f\|_{B_{1,\infty}^{\alpha}(\Omega)}\coloneqq\|f\|_{L_{1}(\Omega)}+|f|_{B_{1,\infty}^{\alpha}(\Omega)}.$
Lastly, $C^{k}(\Omega)$ denotes the set of functions whose coordinate components are $k$-times continuously differentiable with bounded derivatives (in the supremum norm). For a multi-index $k\in\mathbb{N}_{0}^{d}$ with $|k|=$ $\sum_{i=1}^{d}k_{i}$, we write
$D^{k}=\frac{\partial^{|k|}}{\partial x_{1}^{k_{1}}\cdots\partial x_{d}^{k_{d}}},$
where $\frac{\partial}{\partial x_{i}}$ denotes the weak partial derivative. For a function $f\colon\Omega\rightarrow\mathbb{R}$ we denote for $\beta\in\mathbb{N}$
$\|f\|_{C^{\beta}}\coloneqq\max_{k\colon|k|\leq\beta}\|D^{k}f\|_{\infty}.$
For the Jacobian of a vector valued function with respect to the vector $x$ we write $D_{x}$. The covering number of a set $\mathcal{A}$ with respect to the norm $\|\cdot\|$ and the $\tau$-covering is denoted by $\mathcal{N}(\tau,\mathcal{A},\|\cdot\|)$.
2 The Flow Matching objective
@@
Given a fixed latent distribution $\mathbb{U}$ with Lebesgue density $p_{0}$, the vector field $v$ induces a path of probability densities $p_{t}\colon\mathbb{R}^{d}\rightarrow\mathbb{R}_{>0}$, defined via the push-forward:
@@
with $\int p_{t}(x)\mathrm{d}x=1$ for all $t\in[0,1]$.
Using a possibly parameterized function $\tilde{v}$ to approximate a given vector field $v$ results in the Flow Matching objective
$\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ X_{t}\sim p_{t}\end{subarray}}\big{[}|\tilde{v}_{t}(X_{t})-v_{t}(X_{t})|^{2}\big{]}$ (4)
by Lipman et al. (2023). Conversely, in certain cases, a vector field can be determined to match a given density path. Lipman et al. (2023) constructed probability paths resulting from a convolution,
$p_{t}(x)=\int p_{t}(x|y)p^{*}(y)\;\mathrm{d}y,$ (5)
where $p_{t}(\cdot|y)\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a conditional probability path induced through some vector field $v_{t}(\cdot|y)\colon\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ for $y\in\mathbb{R}^{d}$. It can be shown that the vector field inducing (5) is of the form
$v_{t}(x)=\int v_{t}(x|y)\frac{p_{t}(x|y)p^{*}(y)}{p_{t}(x)}dy.$ (6)
In this setting, Lipman et al. (2023) prove that minimizing the Flow Matching objective (4) over a class of parameterized functions $\mathcal{M}$ is equivalent to minimizing the conditional Flow Matching objective given by
$\Psi(\tilde{v})\coloneqq\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big{[}\left|\tilde{v}_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big{]}.$ (7)
As the following result shows, we can calculate the difference between (4) and (7) explicitly.
Lemma 2.1.
In practice, $p^{*}$ is unknown and therefore the empirical counterpart of (7),
$\tilde{\Psi}(\tilde{v})\coloneqq\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ X_{t}\sim p_{t}\left(\cdot|Y_{i}\right)\end{subarray}}\big{[}\left|\tilde{v}_{t}(X_{t})-v_{t}\left(X_{t}|Y_{i}\right)\right|^{2}\big{]}$ (8)
is minimized. Kunkel and Trabs (2025b) show that there is also an equivalence of gradients in the empirical setting. Let $\mathcal{M}$ be a set of parameterized functions that map $\mathbb{R}^{d}$ into $\mathbb{R}^{d}$. From minimizing (8) in $\tilde{v}\in\mathcal{M}$, we obtain an argument $\hat{v}$. Solving the ODE (3) using $\hat{v}$, we obtain a flow $\hat{\psi}_{t}$, i.e. $\hat{\psi}$ is given by
[…]
We use this flow to push forward the known, latent distribution to time $t=1$. In accordance with the goal of generative modeling, the distribution of this pushforward should mimic $\mathbb{P}^{*}.$
3 Wasserstein distance in Flow Matching
The underlying objective of this work is to evaluate the ability of Flow Matching to mimic the unknown distribution $\mathbb{P}^{*}$ depending on the sample size $n$. For this purpose, it is necessary to measure the gap between $\mathbb{P}^{*}$ and $\mathbb{P}^{\hat{\psi}_{1}(Z)}$ with flow $\hat{\psi}$ from (9). We adopt the Wasserstein-$1$ metric for evaluation.
@@ 3 Wasserstein distance in Flow Matching
$\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})$ $\coloneqq\inf_{\pi\in\Pi(\mathbb{P},\mathbb{Q})}\int_{\mathcal{X}}|x-y|\;\mathrm{d}\pi(x,y),$
where $\Pi(\mathbb{P},\mathbb{Q})$ denotes the set of couplings with marginals $\mathbb{P}$ and $\mathbb{Q}$. By Villani (2008, Theorem 5.10(i)), this definition admits the following dual formulation:
$\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})=\sup_{W\in\operatorname{Lip}(1)}\mathbb{E}_{\begin{subarray}{c}X\sim\mathbb{P}\\ Y\sim\mathbb{Q}\end{subarray}}[W(X)-W(Z)].$ (10)
On the space of probability measures with finite first moment, the Wasserstein-$1$ distance characterizes weak convergence. A detailed discussion of its advantages over other metrics with the same characteristic can be found in (Villani, 2008, p. 98 f.). Due to its geometric interpretability and robustness in high-dimensional, manifold-structured data settings, the Wasserstein-$1$ metric has commonly been used to evaluate generative models. Examples include: Schreuder et al. (2021), Liang (2021), Stéphanovitch et al. (2023) or Kunkel and Trabs (2025a).
The following theorem provides a first comparison of the performance of $\mathbb{P}^{\hat{\psi}_{1}(Z)}$ and $\mathbb{P}^{\psi_{1}(Z)}$, where $\psi$ is the solution of the ODE (3) using (6) as the vector field. In order to bound $\mathsf{W}_{1}(\mathbb{P}^{\hat{\psi}_{1}(Z)}\mathbb{P}^{\psi_{1}(Z)})$ in terms of the difference between the underlying vector fields, we use Grönwall’s Lemma as the classical tool to control ODE stability. This leads to the typical exponential bound on the Lipschitz constant in space. The proof follows the along lines of (Albergo and Vanden-Eijnden, 2022, Proposition 3).

### v2

@@
(Karlsruhe Insitute of Technologythanks: The author is now at Ruhr University Bochum.)
@@ 1 Introduction
Flow Matching, introduced by Lipman et al. (2023), has recently attracted attention for its simplicity compared to diffusion models (Sohl-Dickstein et al., 2015), which have long been regarded as state-of-the-art. Unlike generative adversarial networks (Goodfellow et al., 2014), which learn $\psi$ directly, Flow Matching learns a time-depended vector field $v_{t}$ that transports mass from $\mathbb{U}$ to an approximation of $\mathbb{P}^{*}$ via the ODE for $t\in[0,1]$:
@@
Earlier, normalizing flows (Tabak and Vanden-Eijnden, 2010; Tabak and Turner, 2013), popularized by Rezende and Mohamed (2015) and Dinh et al. (2015) used discrete invertible mappings with the change-of-variables formula. Continuous normalizing flows (Chen et al., 2018) extended this to continuous mappings using neural ODEs trained via maximum likelihood, which circumvented the calculation of the determinant of the Jacobian. However, CNFs still require significant computational costs due to the necessity of ODE simulations (Grathwohl et al., 2019). Flow Matching avoids these simulations by using a simple least-squares loss
@@
where $v_{t}(\cdot|Y)$ and $p_{t}(\cdot|Y)$ are known by design. In some settings, this resembles the score matching objective (Hyvärinen, 2005; Vincent, 2011) used in diffusion models (Song et al., 2021). Given $n$ i.i.d. observations from $\mathbb{P}^{*}$, minimizing the empirical counterpart of (2) over a set of functions $\mathcal{M}$ and solving the ODE (1) leads to the estimator $\hat{\psi}$. The statistical question is how well $\mathbb{P}^{\hat{\psi}_{1}(Z)}$ approximates $\mathbb{P}^{*}$ depending on the number of observations $n$ and the function class $\mathcal{M}$.
@@
Algorithms based on Flow Matching have been successfully applied in a range of domains, including text-to-speech (Guo et al., 2024), text-to-image generation (Yang et al., 2025; Esser et al., 2024), molecular and protein structure design (Dunn and Koes, 2024; Bose et al., 2024), and surrogate modeling in high-energy physics (Bieringer et al., 2024). The method has also been extended to various theoretical settings: Atanackovic et al. (2024) adapted it to interacting particle systems, Gat et al. (2024) explored discrete spaces, Chen and Lipman (2024) generalized it to Riemannian manifolds, and Kerrigan et al. (2024) extended it to function spaces.
@@
The analysis of Flow Matching is more fragile compared to diffusions, as Girsanov’s theorem does not apply, thus making the strategy used in Chen et al. (2023c) inapplicable. Using Grönwall’s Lemma, the natural and standard approach for stability estimates of ODEs, results in exponential dependency on the Lipschitz constant of the approximated vector field. Consequently, the statistical results so far are rather limited: Fukumizu et al. (2025) use a slight adaptation and introduce stopping times to transfer some results from diffusion models but they do not circumvent the growth of the Lipschitz constant, which causes difficulties in their proof. Gao et al. (2024b) use the same adaptation and obtain rates in the Wasserstein 2 distance for certain types of unknown distributions. Stéphanovitch et al. (2025) study deterministic ODE sampling, but, due to their general setting, examine the infinite-time setting of diffusions. Kunkel and Trabs (2025b) showed that Flow Matching attains minmax optimal rates, but they use an overparameterized setting which does not align with practical implementations. Due to the excess of parameters, they are able to compensate the rising Lipschitz constant.
@@
In this work, we first focus on a detailed study of the Lipschitz constant of the intrinsic "true" vector field of Flow Matching. By providing upper and lower bounds, we demonstrate that the choice of the variance function and the behavior of the covariance of the reweighted unknown distribution is crucial. This function governs the level of smoothing depending on $t\in[0,1]$. While current results focus on fixed choices of variance functions, we study the Lipschitz constant in terms of this function. We derive assumptions on the unknown distribution that lead to control over the Lipschitz constant for various types of variance functions. Our general assumptions are validated for a class of distributions with unbounded support considered by Stéphanovitch et al. (2025).
@@
Subsequently, we employ Bernstein-type bounds to derive an oracle inequality for estimating the vector field. We carefully adapt these classical methods to address the theoretical challenges of Flow Matching. When bounding the error of the Flow Matching distribution estimator in the Wasserstein $1$ distance, standard approaches from approximation theory usually increase the negative impact of the dimension on the rate. This issue can be circumvented up to a small error by exploiting the smoothness of the vector field. Using this approach allows us to build on existing results on network approximation and thus keep the focus on the distributional aspects of Flow Matching. We consider feedforward rectified linear unit (ReLU) networks, hence we do not require smooth networks for our results. ReLU nets with logarithmically increasing depth and a polynomial number of non-zero weights suffice to achieve improved convergence rates. Additionally, we exploit the density path of the true vector field to take advantage of the smoothness of the unknown distribution. Thus, our rates are faster in high dimensions than those previously obtained in Gao et al. (2024b), while allowing for considerably smaller networks than in Kunkel and Trabs (2025b).
@@
As Lipman et al. (2023) noted, Flow Matching is closely connected to diffusion models. For a comprehensive survey of generative diffusion models, refer to Cao et al. (2024). Even when the model is not built to be consistent with diffusion models, approximating a score function is similar to approximating the vector field in Flow Matching. The statistical analysis of score matching and diffusion models are an active area of research, see for instance Chen et al. (2023a), Chen et al. (2023b), Chen et al. (2023c), Oko et al. (2023), Tang and Yang (2024), Azangulov et al. (2024), Zhang et al. (2024) , Yakovlev and Puchkin (2025), Stéphanovitch et al. (2025).
@@
Section˜2 introduces the problem setting mathematically and recalls all the definitions needed to define the conditional Flow Matching objective from Lipman et al. (2023). The section then discusses the use of the Wasserstein distance and shows a classical first error decomposition using the aforementioned Grönwalls lemma. After that, we introduce the model based on the choice $\mathbb{U}=\mathcal{N}(0,I_{d})$, which will be the focus of the subsequent work. In Section˜3 we establish upper and lower bounds of the Lipschitz constant. We then show that bounding the Lipschitz constant requires careful investigation of the covariance of the reweighted unknown distribution. Next, we demonstrate that, under certain assumptions, we can construct Flow Matching such that the Lipschitz constant is bounded. In Section˜4 we derive an oracle inequality and finally obtain a rate of convergence using feedforward ReLU networks. All proofs are deferred to Section˜6.
@@ 1.1 Some notations
We use the notation $X\sim p$ as shorthand for $X\sim\mathbb{P}$ when the distribution $\mathbb{P}$ of the random variable $X$ admits a density $p$ with respect to the Lebesgue measure. The symbol $\lesssim$ denotes an inequality of the form $a\leq c\cdot b$, where $a,b\in\mathbb{R}$ and $c$ is a constant independent of the sample size $n$. In the same setting, the symbol $a\sim b$ denotes equality up to a constant. For $d\in\mathbb{N}$, we equip $\mathbb{R}^{d}$ with the standard Euclidean norm $|\cdot|$. For a matrix $A\in\mathbb{R}^{d\times d}$, we denote the spectral norm by $\|A\|$. Given a function $f\colon\Omega\rightarrow\mathbb{R}^{d^{\prime\prime}}$, where $\Omega\subset$ $\mathbb{R}^{d^{\prime}}$, the supremum norm and the $L_{1}$-norm are defined by $\|f\|_{\infty,\Omega}\coloneqq\sup_{x\in\Omega}|f(x)|\quad\text{and}\quad\|f\|_{1,\Omega}\coloneqq\int_{\Omega}|f(x)|\;\mathrm{d}x.$ When $\Omega=\mathbb{R}^{d^{\prime}}$, we abbreviate $\|f\|_{\infty,\mathbb{R}^{d^{\prime}}}$ as $\|f\|_{\infty}$, and similarly $\|f\|_{1,\mathbb{R}^{d^{\prime}}}=\|f\|_{1}$. For $d^{\prime\prime}=1$ and $\alpha\in(0,1]$, the Besov space $B_{1,\infty}^{\alpha}(\Omega)$ is defined as $B_{1,\infty}^{\alpha}(\Omega)\coloneqq\big\{f\in L_{1}(\Omega)\colon|f|_{B_{1,\infty}^{\alpha}(\Omega)}\coloneqq\sup_{t>0}t^{-\alpha}\omega_{1}(f,t)_{1}<\infty\big\},$ where the modulus of smoothness is given by $\omega_{1}(f,t)_{1}\coloneqq\sup_{0<|h|\leq t}\int|f(x)-f(x+h)|\;\mathrm{d}x,\;t>0,x\in\Omega.$ If $x+h\notin\Omega$, the integrand is taken to be zero. The norm on $B_{1,\infty}^{\alpha}(\Omega)$ is defined by $\|f\|_{B_{1,\infty}^{\alpha}(\Omega)}\coloneqq\|f\|_{L_{1}(\Omega)}+|f|_{B_{1,\infty}^{\alpha}(\Omega)}.$ Lastly, $C^{k}(\Omega)$ denotes the set of functions whose coordinate components are $k$-times continuously differentiable with bounded derivatives (in the supremum norm). For a multi-index $k\in\mathbb{N}_{0}^{d}$ with $|k|=$ $\sum_{i=1}^{d}k_{i}$, we write $D^{k}=\frac{\partial^{|k|}}{\partial x_{1}^{k_{1}}\cdots\partial x_{d}^{k_{d}}},$ where $\frac{\partial}{\partial x_{i}}$ denotes the weak partial derivative. For a function $f\colon\Omega\rightarrow\mathbb{R}$ we denote for $\beta\in\mathbb{N}$ $\|f\|_{C^{\beta}}\coloneqq\max_{k\colon|k|\leq\beta}\|D^{k}f\|_{\infty}.$ For the Jacobian of a vector valued function with respect to the vector $x$ we write $D_{x}$. The Gradient is denoted by $\nabla$ and the Hessian is denoted by $\nabla^{2}$. The covering number of a set $\mathcal{A}$ with respect to the norm $\|\cdot\|$ and the $\tau$-covering is denoted by $\mathcal{N}(\tau,\mathcal{A},\|\cdot\|)$.
2 The Flow Matching objective and Wasserstein evaluation
@@
Given a fixed latent distribution $\mathbb{U}$ with Lebesgue density $p_{0}$, the vector field $v$ induces a path of probability densities $p\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{>0}$, defined via the push-forward:
@@
with $\int p_{t}(x)\mathrm{d}x=1$ for all $t\in[0,1]$. Lipman et al. (2023) constructed probability paths resulting from a convolution and a corresponding vector field,
$p_{t}(x)=\int p_{t}(x|y)p^{*}(y)\;\mathrm{d}y,\quad v_{t}(x)=\int v_{t}(x|y)\frac{p_{t}(x|y)p^{*}(y)}{p_{t}(x)}dy.$ (4)
where $p_{t}(\cdot|y)\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a conditional probability path induced through some vector field $v_{t}(\cdot|y)\colon\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ for $y\in\mathbb{R}^{d}$. It can be shown that $v$ induces $p$ in (4). In order to approximate $v$ with a parameterized function $\tilde{v}$, which is a neural network in practice, Lipman et al. (2023) consider the following objective
$\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ X_{t}\sim p_{t}\end{subarray}}\big[|\tilde{v}_{t}(X_{t})-v_{t}(X_{t})|^{2}\big]=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big[\left|\tilde{v}_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big]+C,$ (5)
where $C$ is a constant that is explicitly calculated in Lemma˜A.1. In practice, $p^{*}$ is unknown and therefore the empirical counterpart of (5) in minimized over some class of parameterized function $\mathcal{M}$. Kunkel and Trabs (2025b) show that there is also an equivalence of gradients in the empirical setting. The estimated flow $\hat{\psi}$ is obtained by
$\frac{d}{dt}\hat{\psi}_{t}(x)=\hat{v}_{t}(\hat{\psi}_{t}(x)),\quad\hat{\psi}_{0}(x)=x,\qquad\text{for}\quad\hat{v}\in\operatorname*{argmin}_{\tilde{v}\in\mathcal{M}}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ X_{t}\sim p_{t}\left(\cdot|Y_{i}\right)\end{subarray}}\big[\left|\tilde{v}_{t}(X_{t})-v_{t}\left(X_{t}|Y_{i}\right)\right|^{2}\big].$ (6)
We use this flow to push forward the known, latent distribution to time $t=1$. In accordance with the goal of generative modeling, the distribution of this pushforward, $\mathbb{P}^{\psi_{1}(Z)}$, should mimic $\mathbb{P}^{*}.$
For this purpose, it is necessary to measure the gap between $\mathbb{P}^{*}$ and $\mathbb{P}^{\hat{\psi}_{1}(Z)}$ with flow $\hat{\psi}$ from (6). We adopt the Wasserstein-$1$ metric for evaluation.
@@
$\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})\coloneqq\inf_{\pi\in\Pi(\mathbb{P},\mathbb{Q})}\int_{\mathcal{X}}|x-y|\;\mathrm{d}\pi(x,y)=\sup_{W\in\operatorname{Lip}(1)}\mathbb{E}_{\begin{subarray}{c}X\sim\mathbb{P}\\ Y\sim\mathbb{Q}\end{subarray}}[W(X)-W(Z)],$ (7)
where $\Pi(\mathbb{P},\mathbb{Q})$ denotes the set of couplings with marginals $\mathbb{P}$ and $\mathbb{Q}$ and the equality is shown in Villani (2008, Theorem 5.10(i)). On the space of probability measures with finite first moment, the Wasserstein-$1$ distance characterizes weak convergence. A detailed discussion of its advantages over other metrics with the same characteristic can be found in Villani (2008, p. 98 f.). Due to its geometric interpretability and robustness in high-dimensional, manifold-structured data settings, the Wasserstein-$1$ metric has commonly been used to evaluate generative models. Examples include: Schreuder et al. (2021), Liang (2021), Stéphanovitch et al. (2024) or Kunkel and Trabs (2025a).
The following theorem provides a first comparison of the performance of $\mathbb{P}^{\hat{\psi}_{1}(Z)}$ and $\mathbb{P}^{\psi_{1}(Z)}$, where $\psi$ is the solution of the ODE (3) using (4) as the vector field. In order to bound $\mathsf{W}_{1}(\mathbb{P}^{\hat{\psi}_{1}(Z)},\mathbb{P}^{\psi_{1}(Z)})$ in terms of the difference between the underlying vector fields, we use Grönwalls Lemma as the classical tool to control ODE stability. This leads to the typical exponential bound on the Lipschitz constant in space. The proof follows the along lines of Proposition 3 in Albergo and Vanden-Eijnden (2023).